OpenStudy (anonymous):

Write an algebraic equation for the following: The quotient of a number q and the quantity four times r is equal to seven

OpenStudy (anonymous):

how do you write four times r?

OpenStudy (mathmale):

In other words, ctthomas, how do you express "four times r" algebraically?

OpenStudy (mathmale):

ctt: do you want to solve this problem, or are you more interested in helping others with their problems (as you seem to be doing now)?

OpenStudy (anonymous):

I want to solve it. That's how they stated the question.

OpenStudy (mathmale):

OK, welcome back. Please see whether you can follow Amit's suggestion, above.

OpenStudy (anonymous):

teltoamit 4r

OpenStudy (anonymous):

add this to q, that equals to7

OpenStudy (anonymous):

4r+q=7 ??

OpenStudy (anonymous):

doesn't quotient mean divide? So it would be q/4r=7

OpenStudy (mathmale):

CTT: You did well. Your result is correct. But it'd be clearer if you'd please enclose that 4r inside parentheses: q/(4r) = 7.

OpenStudy (anonymous):

mathmale thank you!!:)